An archaeologist numbers some articles 1 to 11 in the order he discovers them. He selects at random a sample of 3 of them. What is the probability that the sum of the numbers on the items he selects is less than or equal to 8? (You do not need to list all combinations of 3 items from 11 to answer this question.)  If the archaeologist believed that items belonging to the more recent of two civilizations were more likely to be found earlier in his dig and of his 11 items 3 are identified as belonging to that more recent civilization (but the remaining 8 come from an earlier civilization) does a rank sum of 8 for the 3 matching the more recent civilization provide reasonable support for his theory?
Answer #1

Solution

Back-up Theory

Number of ways of selecting r things out of n things is given by nCr = (n!)/{(r!)(n - r)!}…..................................................…(1)

Values of nCr can be directly obtained using Excel Function: Math & Trig COMBIN(Number, Number_chosen)

[Number is n, Number_chosen is r]………………………….............................................................................……………. (1a)

Probability of an event E, denoted by P(E) = n/N ………...................................................……………………………………(2)

where n = n(E) = Number of outcomes/cases/possibilities favourable to the event E and N = n(S) = Total number all possible outcomes/cases/possibilities.

Now to work out the solution,

First we note that out of 11 articles, a random a sample of 3 articles can be selected in

11C3 = (11!)/{(3!)(8)!} [vide (1)]

= (11 x 10 x 9)/(6)

= 165 ways.

And so, vide (2), N = 165 ………………………………………………………........................................................…………….(3)

Part (a)

We note that the order in which the three numbers appear in the sample is of no consequence since matter of concern is only the sum.

We also note that the three numbers in the sample must be distinct.

A sum of 8 or less => 8, 7 or 6 [note that the minimum sum three distinct integers is 6]   

There are only 4 possible combinations which would yield the above sums, namely

(1, 2, 5), (1, 3, 4), (1, 2, 4) and (1, 2, 3).

Thus, vide (2), n = 4 and hence the required probability = 4/165

= 0.0242 Answer

Part (b)

Since the archaeologist has numbered the articles 1 to 11 in the order he discovered them, it is fair to infer that articles numbered 6 to 11 [i.e., half the number of articles] pertain to the recent civilization and articles numbered 1 to 5 pertain the earlier civilization.

Three distinct numbers greater than 5 cannot give a sum of 8. Thus, a rank sum of 8 or less does not provide reasonable support for his theory that items belonging to the more recent of two civilizations were more likely to be found in the sample. Answer

DONE